Please help me with this problem.
What number is 15 and 1/2 percent of 1,200?
please work out this problem for me to show how you got this problem.
Thank you.

Answer by tutorcecilia(2144)

Use the formula:
Part/Whole = %/100
.
Plug-in the known values. Plug-in "x" for the unknown value.
x/1200 = 15.5/100
.
.
Cross multiply:
x/1200 = 15.5/100
.
.
100x = (1200)(15.5)
.
.
Solve for x:
x = (1200)(15.5)/100
x= 186
Check the answer:
186/1200 = 15.5/100